### 23 Dec 2023 v5

@@ Abstract
journal: Journal of Nonlinear Science
@@ 4.1 Psychophysically plausible parameters for a Wilson-Cowan model in V1
For the wavelet representation here we assume 4-orientation steerable transforms Simoncelli92 as a convenient model of the simple cells (as done in Schwartz01 ; Martinez17 ; Martinez19 ). In the experiments involving the (computationally intensive) integration of the Wilson-Cowan differential equation, Section 4.2, we used wavelets with 3 scales in $40\times 40$ images to speed up the computation. But in the psychophysical illustrations, Section 4.3, we used 4 scales in $64\times 64$ images.
@@ 5 Final remarks
Finally, the relation between models proposed here opens the possibility to analyze Divisive Normalization from new perspectives, following methods that have been developed for Wilson-Cowan systems Sejnowski09 . Similarly, mechanisms that generalize the Wilson-Cowan equation such as the neurons with intrinsically nonlinear receptive fields Bertalmio20 could be analyzed via information theoretic tools that have been used to quantify the performance of Divisive Normalization Malo20 ; Malo22 ; Serences14 .
@@ Appendix A Activation functions
Figure 4 in the main text shows that both functions share the same qualitative properties: they are saturating sigmoids, with derivatives that peak at the origin and decrease with the signal.

### 28 Dec 2023 v6

@@ Abstract
journal: J. Nonlinear Science
@@ 4.1 Psychophysically plausible parameters for a Wilson-Cowan model in V1
For the wavelet representation here we assume 4-orientation steerable transforms Simoncelli92 as a convenient model of the simple cells (as done in Schwartz01 ; Martinez17 ; Martinez19 ). In the experiments involving the (computationally intensive) integration of the Wilson-Cowan differential equation, Section 4.2, we used wavelets with 3 scales in $40\times 40$ images to speed up the computation. But in the psychophysical illustrations, Section 4.3, we used 4 scales in $64\times 64$ images. Of course these choices have to be considered as illustrative options. The selected model Martinez17 ; Martinez19 , has the elements of previous image-computable models Watson97 ; Winawer13 ; Schutt17 , but of course, the different instances differ in details despite they have the same qualitative flavour.
@@ 5 Final remarks
Finally, the relation between models proposed here opens the possibility to analyze Divisive Normalization from new perspectives, following methods that have been developed for Wilson-Cowan systems Sejnowski09 . Similarly, mechanisms that generalize the Wilson-Cowan equation such as the neurons with intrinsically nonlinear receptive fields Bertalmio20 could be analyzed via information theoretic tools used to quantify the transmission performance of Divisive Normalization Serences14 ; Malo20 ; Malo22 , or the functional connectivity Li_24 .
@@ Appendix A Activation functions
Figure 4 in the main text shows that both functions share the same qualitative properties: saturating sigmoids with derivatives that peak at the origin and decrease with the signal.
@@ Appendix B Matlab code
Appendix C Small-scale Divisive Normalization
Overview.
Similarly to Gomez19 the code includes a reduced-scale model which can be applied to 3-pixel images for full visualization of the responses and the phase-space. This reduced-scale model consists of two linear+nonlinear layers: (1) a linear radiance-to-luminance transform using a standard Spectral Sensitivity Function, $V_{\lambda}$, in the spectral integration Stiles82 , followed by a simple exponential for the luminance-to-brightness nonliniearity applied pixel-wise in the spatial domain, that simulates the Weber-Fechner response to luminance Fairchild13 , and (2) a linear+nonlinear layer in which the linear transform is a discrete cosine transform (a orthonormal rotation) followed by a low-pass weighting function that simulate frequency-tuned sensors and the Contrast Sensitivity Function (CSF) Campbell68 . Then, the outputs of the frequency sensors undergo a nonlinear interaction that may be a Divisive Normalization Carandini94 ; Carandini12 ; Martinez17 ; Martinez19 , or its equivalent Wilson-Cowan network, with parameters computed according to Eq. 13.
${\boldsymbol{x}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
$\mathcal{L}^{(1)}$
${\boldsymbol{r}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
$\mathcal{N}^{(1)}$
${\boldsymbol{x}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
$\mathcal{L}^{(2)}$
${\boldsymbol{r}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
$\mathcal{N}^{(2)}$
${\boldsymbol{x}^{2}}$
Transform.
The actual inputs of our code are the responses of the linear photoreceptors: 3-pixel image vectors with normalized luminance values, i.e. $\boldsymbol{r}^{1}\in\mathbb{R}^{3}$. The normalized luminance was computing dividing the absolute luminance in $cd/m^{2}$ by the value corresponding to the 95% percentile of the luminance, in our case 260 $cd/m^{2}$.
  • The luminance-to-brightness transform, $\mathcal{N}^{(1)}$, is just:
    $\boldsymbol{x}^{1}=(\boldsymbol{r}^{1})^{\gamma}\,\,\,\,\,\,\,\,\,\,\textrm{where}\,\,\,\,\gamma=0.6$ (21)
  • The linear transform of frequency-tuned sensors with CSF gain, $\mathcal{L}^{(2)}$, is:
    $\boldsymbol{r}^{2}=G_{\textrm{CSF}}\cdot F\cdot\boldsymbol{x}^{1}\,\,\,\,\,\,\,\,\,\,\textrm{where}$ (22)
    $F$ $=$ $\left(\begin{array}[]{ccc}\sqrt{\frac{1}{3}}&\sqrt{\frac{1}{3}}&\sqrt{\frac{1}{3}}\\[8.5359pt] \sqrt{\frac{1}{2}}&0&-\sqrt{\frac{1}{2}}\\[8.5359pt] -\sqrt{\frac{1}{6}}&\sqrt{\frac{2}{3}}&-\sqrt{\frac{1}{6}}\\ \end{array}\right)$
[…]
  • The Divisive Normalization of the frequency-tuned sensors, $\mathcal{N}^{(2)}_{\textrm{DN}}$, is:
    $\boldsymbol{x}^{2}=sign(\boldsymbol{x}^{2})\odot\mathbb{D}_{\boldsymbol{k}}\cdot\mathbb{D}^{-1}_{\left(\boldsymbol{b}+\boldsymbol{H}\cdot|\boldsymbol{r}^{2}|^{\gamma}\right)}\cdot|\boldsymbol{r}^{2}|^{\gamma}\,\,\,\,\,\,\textrm{where}\,\,\,\,\,\,\gamma=0.7,\,\,\,\,\,\,\textrm{and,}$, (25)
[…]
    and the vector of semisaturations, $\boldsymbol{b}$, is:
    $\boldsymbol{b}=\left(\begin{array}[]{c}0.08\\ 0.03\\ 0.01\\ \end{array}\right)$
  • The equivalent Wilson-Cowan interaction, $\mathcal{N}^{(2)}_{\textrm{WC}}$, is defined by the differential equation 8, where the auto-attenuation, $\boldsymbol{\alpha}$, and the interaction matrix, $\boldsymbol{W}$, are:
    $\boldsymbol{\alpha}$ $=$ $\left(\begin{array}[]{c}0.41\\ 1.10\\ 1.30\\ \end{array}\right)$ (31)
[…]
    and the saturation function is:
    $f(\boldsymbol{x})=c\,\boldsymbol{x}^{\gamma}\,\,\,\,\,\,\textrm{where}\,\,\,\,\,\,\gamma=0.4,\,\,\,\,\,\,\textrm{and,}$, (36)
    the scaling constant is, $c=\hat{\boldsymbol{x}}^{1-\gamma}$, and $\hat{\boldsymbol{x}}$ is the average response over natural images (for the Divisive Normalization transform):
    $\hat{\boldsymbol{x}}=\left(\begin{array}[]{c}1.12\\ 0.02\\ 0.01\\ \end{array}\right)$
    This exponent is also used for the definition of energy in Wilson-Cowan, $\boldsymbol{e}=|\boldsymbol{r}|^{\gamma}$.
Note that the interaction neighborhoods have unit volume, $\sum_{j}W_{ij}=1\,\,\,\forall j$, as suggested in Watson97 , and then, the Divisive Normalization kernel is given by the product of this unit-volume neighborhood and two left and right filters in the diagonal matrices, $\mathbb{D}_{l}$ and $\mathbb{D}_{r}$ Martinez19 . The values for the semisaturation, $\boldsymbol{b}$, and the diagonal matrices $\mathbb{D}_{l}$ and $\mathbb{D}_{r}$ were inspired by the contrast response results in Martinez19 : we set the semisaturation according to the average response of natural images (low-pass in nature), and we initialized the left and right filters to high-pass. However, afterwards, in order to make $\mathcal{N}_{\textrm{DN}}$ and $\mathcal{N}_{\textrm{WC}}$ consistent, we applied the Divisive Normalization over natural images and we iteratively updated the values of the right and left filters according to Eq. 13. In the end, we arrived to the values in the above expressions (where the filter at the left is high-pass, but the filter at the right is not). Note that the attenuation in Wilson-Cowan is computed using Eq. 13.
Jacobian.
The perceptual metric Martinez17 and information transmission Gomez19 ; Malo20 strongly depend on how the system (locally) deforms the signal representation. This is described by the Jacobian of the transform with regard to the signal, $\nabla_{\boldsymbol{r}^{1}}S=\nabla_{\boldsymbol{r}^{2}}\mathcal{N}^{(2)}\cdot\nabla_{\boldsymbol{x}^{1}}\mathcal{L}^{(2)}\cdot\nabla_{\boldsymbol{r}^{1}}\mathcal{N}^{(1)}$. In this reduced-scale model, this Jacobian (for the Wilson-Cowan case) is:
$\nabla_{\boldsymbol{r}^{1}}S=\gamma_{2}\left(\mathbb{D}_{\boldsymbol{\alpha}}+\boldsymbol{W}\cdot\mathbb{D}_{\frac{df}{dx}}\right)^{-1}\cdot\left(\mathbb{D}_{\left(\mathbb{D}_{\boldsymbol{\alpha}}\cdot\boldsymbol{x}^{2}+\boldsymbol{W}\cdot f(\boldsymbol{x}^{2})\right)}\right)^{1-\frac{1}{\gamma_{2}}}\cdot G_{\textrm{CSF}}\cdot F\cdot\mathbb{D}_{\gamma_{1}(\boldsymbol{r}^{1})^{\gamma_{1}-1}}$ (37)